OpenStudy (anonymous):

PLZ HELP! Medal given!! Simplify: mn^-4/p^0q^-2

OpenStudy (anonymous):

The way its worded is weird... So here is a pic of the equation.

OpenStudy (aaronq):

\(\dfrac{1}{q^{-2}}=q^2\), any number to the power of zero is 1 so \(p^0=1\) so \(\dfrac{mn^{-4}}{p^0q^{-2}}=\dfrac{mq^{2}}{n^{4}}\)